OpenStudy (anonymous):

Evaluate the indefinite integral:

OpenStudy (anonymous):

\[\int\limits_{}^{}(\ln(x))^9/x)dx \]

OpenStudy (anonymous):

let u = ln(x), so then du = 1/x. Then your integral becomes \[\int\limits_{?}^{?}u^9du\] = \[(1/10)(\ln(x))^(10) +c\]. That's ln to the tenth just in case you're confused.